### v2

# Towards Interactive and Learnable Cooperative Driving Automation: a Large Language Model-Driven Decision-Making Framework

Abstract
At present, Connected Autonomous Vehicles (CAVs) have begun to open road testing around the world, but their safety and efficiency performance in complex scenarios is still not satisfactory. Cooperative driving leverages the connectivity ability of CAVs to achieve synergies greater than the sum of their parts, making it a promising approach to improving CAV performance in complex scenarios. However, the lack of interaction and continuous learning ability limits current cooperative driving to single-scenario applications and specific Cooperative Driving Automation (CDA). To address these challenges, this paper proposes CoDrivingLLM, an interactive and learnable LLM-driven cooperative driving framework, to achieve all-scenario and all-CDA. First, since Large Language Models(LLMs) are not adept at handling mathematical calculations, an environment module is introduced to update vehicle positions based on semantic decisions, thus avoiding potential errors from direct LLM control of vehicle positions. Second, based on the four levels of CDA defined by the SAE J3216 standard, we propose a Chain-of-Thought (COT) based reasoning module that includes state perception, intent sharing, negotiation, and decision-making, enhancing the stability of LLMs in multi-step reasoning tasks. Centralized conflict resolution is then managed through a conflict coordinator in the reasoning process. Finally, by introducing a memory module and employing retrieval-augmented generation, CAVs are endowed with the ability to learn from their past experiences. We validate the proposed CoDrivingLLM through ablation experiments on the negotiation module, reasoning with different shots experience, and comparison with other cooperative driving methods. The results demonstrate that our method provides significant safety and efficiency advantages, as well as the ability to interact and learn effectively in complex environments. Our code is available at: https://github.com/FanGShiYuu/CoDrivingLLM.
Index Terms:
connected autonomous vehicles, cooperative driving automation, large language model, conflict negotiation, retrieval augment generation
I Introduction
As autonomous driving technology continues to advance, we are entering an era where both Connected Autonomous Vehicles (CAVs) and Human-Driven Vehicles (HDVs) will coexist. While CAVs are considered to have great potential in improving traffic safety and efficiency, their current performance on open roads is far from satisfactory. According to California’s Department of Motor Vehicles [1], 51% of disengagements were due to CAVs’ decision-making failures. Additionally, the Beijing Autonomous Vehicle Road Test Report revealed that up to 91% of disengagements occurred during interactions with other vehicles, indicating that current autonomous driving technology is not yet adequate for complex interaction scenarios [2].
To ameliorate this issue, a promising approach is to leverage the cooperative driving capabilities of CAVs [3]. The SAE J3216 standard divides Cooperative Driving Automation (CDA) into four levels: Status-sharing, Intent-sharing, Agreement-seeking, and Prescriptive [4]. Federal Highway Administration has conducted the CARMA Program to develop and test CDA. However, the current validation of the program primarily focuses on single CDA under single scenarios, such as cooperative platooning at highway [5]. How to achieve all-scenario and all-CDA cooperative driving still remains to be tackled.
Various approaches have been proposed to address the problem of cooperative driving in different scenarios. The currently popular cooperative driving methods can be broadly categorized into optimization-based, rule-based, and machine learning-based approaches. Optimization-based methods aim to maximize or minimize an objective function to achieve specific goals. Due to the complexity of cooperation involving multiple participants and relationships, optimization-based research often introduces additional simplifications to balance computational time and performance. A common practice in intersection management is to divide the area into multiple zones, with each zone being occupied by only one vehicle at any given time[6, 7]. An alternative approach is to decompose the problem into two layers and optimize the upper and lower layers respectively to simplify the complexity of the problem [8]. However, optimization-based approaches often fail to explicitly consider regulations and social norms during driving, making their decision results difficult to understand. In contrast, the rule-based approach boasts simplicity in form and thus computational efficiency [9]. Nevertheless, despite the rule-based approaches that can be coupled with traffic regulations, pre-set rules often result in poor robustness. On the other hand, machine learning methods, such as deep learning and reinforcement learning, have gained increasing popularity in recent years. These methods have been successfully applied to cooperative driving in environments such as intersections [10, 11, 12], merging zones [13, 14], and highways [15], demonstrating good performance. However, the performance of these learning-based models often degrades significantly when transferred to untrained environments [16, 17]. Overall, although many studies have made numerous attempts at cooperative driving, the significant differences in driving rules and scene characteristics across various scenarios have led to most research focusing on single scenarios. To achieve all-scenario and all-CDA, it is essential to establish a cooperative driving method that possesses both interactive and learnable capabilities.
On the other hand, the rapid development of Transformer models and Large Language Models (LLMs) in recent years has provided a new potential avenue for achieving cooperative decision-making. These models have already shown significant potential in fields such as natural language processing and decision-making for agents. Sha et al. [18] leveraged LLMs’ common sense and high-dimensional information understanding capabilities to handle rare situations, such as a malfunctioning vehicle stalled in the middle of a roundabout. Compared to learning-based and optimization-based single-vehicle decision-making methods, the overall cost was reduced by 18.1% and 16.4%, respectively. Furthermore, to improve the interaction ability of CAV, research has introduced human instructions to establish LLM decision models aligned with human logic and results showed that Human-in-the-loop guidance for LLMs has significantly improved the interpretability of decisions [19, 20]. Building on this, Cui et al. [21] proposed the Talk2Drive framework, the first-of-its-kind multi-scenario field experiment deploying LLMs on a real-world autonomous vehicle. Experimental results showed that Talk2Drive significantly reduced takeover rates in scenarios such as highways, intersections, and parking lots, with at least a 56.6% reduction in takeover rates. In light of the poor performance of current autonomous driving systems in facing long-tail corner cases, Fu et al. [22] proposed an LLM decision-making framework that includes three main functions: reasoning, interpretation, and memorization. This framework enables CAV to dynamically learn through the interaction with the environment, thereby improving its decision-making ability. On this basis, Wen et al. [23] established the Dilu framework, further introducing a reflection module to identify safe and unsafe decisions produced by the reasoning module, subsequently refining them into correct decisions using the knowledge embedded in the LLM, thus further enhancing the learning ability of CAVs. Although LLMs have shown great potential in autonomous driving decision-making tasks, current research is still primarily focused on single-vehicle decision tasks. Designing an efficient, safe, and socially compatible continuous learning framework based on LLMs for multi-vehicle cooperative driving tasks remains a significant area of exploration.
Against this backdrop, given the significance of cooperative driving in addressing the current challenges faced by CAV in complex scenarios and the potential of LLM in enhancing the interaction and learning ability of CAV, this paper proposes an interactive and learnable LLM-driven cooperative driving framework for all-scenario and all-CDA. The contributions of this paper are presented as follows:
Our contributions can be summarized as follows:
  • This paper proposes CoDrivingLLM, a LLM-based interactive and learnable cooperative driving framework, which features a centralized-distributed coupled architecture. Through the effective design and coupling of the environment module, reasoning module, and memory module, the framework significantly enhances the performance of cooperative driving under different scenarios.
  • In the reasoning module, four sub-modules including state-perception, intent-sharing, negotiation, and decision are designed to enable flexible switching between different levels of CDA. Considering the high complexity of multi-vehicle cooperation, a conflict coordinator that infers the suggested passing order for vehicles is designed as the negotiation module. Results show that the introduction of the conflict coordinator significantly improves the vehicle interaction ability.
  • By introducing the memory module with Retrieval Augment Generation (RAG), CAVs are endowed with the ability to continuously learn from historical experiences. By searching for memories most similar to the current scenario and referring to the corresponding decisions, CAVs can avoid repeating past mistakes. Results demonstrated excellent performance across various scenarios and tasks.
The rest of the paper is organized as follows. Section II formulates the problem of our research. Section III introduces the architecture of the proposed framework. In Section IV, we validate our framework through several experiments. Finally, conclusions are made in Section V.
II Problem Formulation
With the increasing maturity of autonomous driving technology, manufacturers have shifted from focusing on early-stage technical competition to prioritizing commercial deployment. However, CAVs still exhibit various problems on open roads and even become the culprit of many congestion or accidents. In certain scenarios, the accident rate of CAVs even reached 5.25 times that of human drivers, which has failed to meet people’s expectations for autonomous driving technology, and is gradually eroding people’s trust in CAVs. The communication capabilities of CAVs enable connectivity and mutual assistance. Therefore, leveraging cooperative driving capabilities is a promising way to enhance CAV performance.
Additionally, the cooperative decision-making problem for multiple CAVs can be modeled as a partially observable Markov decision process (POMDP)[24]. We define the POMDP using the tuple $\mathcal{M}_{\mathcal{G}}=(\mathcal{V},S,[\mathcal{O}_{i}],[\mathcal{A}_{i}],\mathcal{P},[r_{i}])$, where $\mathcal{V}$ represents the finite set of all controlled agents (CAVs), and $S$ denotes the state space encompassing all agents. $\mathcal{O}_{i}$ represents the observation space for each agent $i\in\mathcal{V}$, $\mathcal{A}_{i}$ denotes the action space, and $r_{i}$ is the reward associated with CAV $i$. The transition distribution is represented by $\mathcal{P}$.
At any given time, each agent $i$ receives an individual observation $o_{i}:\mathcal{S}\to\mathcal{O}_{i}$ and selects an action $a_{i}\in\mathcal{A}_{i}$ based on a policy $\pi_{i}:\mathcal{O}_{i}\times\mathcal{A}_{i}\to[0,1]$. The agent then transitions to a new state $s_{i}^{\prime}$ with a probability given by $\mathcal{P}(s^{\prime}|s,a):\mathcal{S}\times\mathcal{A}_{1}\times...\times\mathcal{A}_{N}\to\mathcal{S}$.
II-1 Observation Space
Due to the limitations of sensor hardware, a CAV can only detect the status information of surrounding vehicles within a limited distance $\mathcal{L}$. We denote the set of all observable vehicles within the perception range of agent $i$ as $\mathcal{N}_{i}$. The observation matrix for agent $i$, denoted as $\mathcal{O}_{i}$, is a matrix with dimensions $|\mathcal{N}_{i}|\times|\mathcal{F}|$, where $|\mathcal{N}_{i}|$ represents the number of observable vehicles for agent $i$, and $|\mathcal{F}|$ represents the number of features used to describe a vehicle’s state. The feature vector for vehicle $k$ is expressed as
$\mathcal{F}_{k}=[x_{k},y_{k},v^{x}_{k},v^{y}_{k},\cos{\phi_{k}},\sin{\phi_{k}}],$ (1)
where $x_{k}$ and $y_{k}$ are the longitudinal and lateral positions, $v^{x}_{k}$ and $v^{y}_{k}$ are the longitudinal and lateral speeds, and $\cos{\phi_{k}}$ and $\sin{\phi_{k}}$ are the cosine and sine of the vehicle’s heading angle $\phi_{k}$, respectively. The overall observation space of the system is the combined observation of all CAVs, i.e., $\mathcal{O}=\mathcal{O}_{1}\times\mathcal{O}_{2}\times\cdots\times\mathcal{O}_{|\mathcal{V}|}$.
II-2 Action Space
[…]
Figure 1: Cooperative Driving with LLM.
III Interactive and learnable LLM-driven cooperative driving framework
Leveraging the cooperative driving ability of CAV can significantly improve the safety and efficiency of traffic systems. However, the current research often focuses on single-scenario, single-function cooperation, leading to insufficient interaction capabilities and a lack of continuous learning capabilities. Utilizing the extensive world knowledge and strong reasoning capabilities of LLMs holds promise for addressing these challenges and achieving all-scenario and all-level cooperative automation, thereby fully unleashing the potential of CAVs to transform traffic systems. As a solution, we propose CoDrivingLLM: an interactive and learnable LLM-driven cooperative driving framework for all scenarios and all cooperative driving automation.
CoDrivingLLM mainly contains three modules: the environment module, the reasoning module, and the memory module. In this section, we first introduce the overall framework of CoDrivingLLM and then provide detailed explanations of each module.
III-A Overall Architecture
Fig. 2 illustrates the main modules and their logical relationships within CoDrivingLLM, which consists of three primary modules: the environment module, the reasoning module, and the memory module. First, the environment module updates the current scene information, including the state of all vehicles such as position, speed, etc., based on the actions of both CAVs and HDVs from the previous time step. Next, we designed a centralized-distributed coupled LLM reasoning module. Based on the four levels of CDA defined by the SAE J3216 standard, we integrated four sub-functions into this reasoning module: state sharing, intent sharing, negotiation, and decision. By incorporating the Chain-of-Thought (COT) method, we sequentially connect each sub-function in the reasoning process to enhance the safety and reliability of decision-making. During this process, each CAV performs distributed high-level logical reasoning using LLMs, completing cooperative driving at different levels, and utilizes the conflict coordinator within the framework for centralized conflict resolution, further enhancing safety. Finally, the scenario description, conflict description, and final decisions from the reasoning process are stored in a memory database with a vectorized form. In the subsequent reasoning, the CAV can reference the most similar past memory as experience, enabling the designed CAV to continuously learn and improve its capabilities while driving.
Figure 2: The overview of CoDrivingLLM, an Interactive and Learnable LLM-Driven Cooperative Driving Framework.
III-B Environment Module
The environment module consists of two sub-modules: (1) the environmental dynamics simulation sub-module, which simulates real-world environmental dynamics to provide realistic background traffic flow and training feedback for the cooperative driving framework, and (2) the model-based control execution sub-module, which provides model-based vehicle control units for the LLM, enhancing the accuracy and success rate of action execution.
III-B1 Environmental Dynamics Simulation Sub-module
Real-world feedback is crucial for training a stable and reliable cooperative decision-making model. To ensure the realism and reliability of the simulation, we consider a mixed human-autonomous driving environment within the environmental dynamics module, introducing uncontrolled human-driven vehicles to create a more realistic background traffic flow for the cooperative driving framework. These uncontrolled vehicles do not participate in cooperative driving tasks, meaning they do not share their intent and operate solely based on their own decision logic.
Given that the combination of IDM (Intelligent Driver Model) and MOBIL (Minimizing Overall Braking Induced by Lane changes) is widely used to characterize human driving behavior and has shown good results in various scenarios such as intersections, roundabouts, and merge areas, we utilize IDM and MOBIL to represent the longitudinal and lateral behaviors of HDVs, respectively.
The acceleration of the HDV is given by the IDM as shown below:
$a_{f}(s,v,\Delta v)=a_{\max}\left[1-\left(\frac{v}{v_{\mathrm{d}}}\right)^{\delta}-\left(\frac{s^{*}(v,\Delta v)}{s}\right)^{2}\right]$ (2)
$s^{*}(v,\Delta v)=s_{0}+vT_{g}+\frac{v\Delta v}{2\sqrt{a_{\max}a_{dd}}}$
where $a_{f}(s,v,\Delta v)$ is the acceleration derived from IDM, $v_{\mathrm{d}}$ is the desired velocity, $\delta$ is the acceleration exponent, $\Delta v$ and $s$ are the velocity difference and distance between the front vehicle and the subject vehicle, $s^{*}(v,\Delta v)$ is the expected distance, $s_{0}$ is the minimum stopping distance, $T_{g}$ is the desired time gap, $a_{\max}$ is the maximum acceleration, and $a_{dd}$ is the desired deceleration.
Furthermore, MOBIL achieves safe and efficient traffic flow by minimizing the overall braking caused by lane change, which mainly includes two parts: lane change incentive and safety inspection. The lane change incentive evaluates the change in the acceleration of the lane change on itself and surrounding vehicles to determine whether a lane change is warranted.
$\underbrace{\tilde{a}_{c}-a_{c}}_{\text{ego-vehicle}}+p(\underbrace{\tilde{a}_{n}-a_{n}}_{\text{new follower}}+\underbrace{\tilde{a}_{o}-a_{o}}_{\text{old follower}})\geq\Delta a_{\text{th}},$ (3)
where $\tilde{a}$ and $a$ are the acceleration of the vehicles before and after the lane change, respectively, $p$ is a politeness coefficient that determines the degree to which the vehicle pays attention to the acceleration changes of the surrounding vehicles when making lane change decisions, and $a_{\text{th}}$ the acceleration gain required to trigger a lane change.
At the same time, in order to ensure the safety of lane change, the MOBIL model also carries out safety inspection to ensure that lane change will not cause a sudden brake on the rear vehicle of the expected lane, that is:
$\tilde{a}_{n}\geq-b_{safe},$ (4)
where $b_{safe}$ is the maximum braking imposed on a vehicle during a cut-in.
III-B2 Model-based Control Execution Sub-module
While LLMs possess strong reasoning capabilities, they do not perform well in precise mathematical calculations and low-level vehicle motion control. To address this, we designed a model-based control execution sub-module. In this model, we use a model-based control approach to translate the semantic outputs of the reasoning module into acceleration and front wheel angles, which are then used to update the vehicle’s position, speed, and other information at the next time step. Specifically, both acceleration and front wheel angles are controlled by relatively simple proportional controllers, with the calculation formulas as follows:
$\begin{split}a&=K_{p}(v_{r}-v)=K_{a}a_{r}\Delta t\\ \delta&=\text{arcsin}(\frac{K_{h}(\phi_{r}-\phi)l}{2v})\end{split}$ (5)
where $K_{a}$ and $K_{h}$ are the proportional longitudinal control and control gain, $v_{e}$ and $a_{r}$ are the expect velocity and reference acceleration, $\phi_{r}$ is reference heading to follow the lane, and $\phi$ is the heading.
After obtaining the vehicle’s acceleration and front wheel angle from the control module, a bicycle model is selected as the kinematic module to generate the states of all vehicles in the environment at the next time step. The specifics are as follows:
$f(a,\delta)=\left\{\begin{aligned} &x=v\cos(\phi+\beta)\\ &y=v\sin(\phi+\beta)\\ &v=a\\ &\phi=\frac{v}{l}\sin\beta\\ &\beta=\tan^{-1}(\frac{1}{2\tan\delta}),\end{aligned}\right.$ (6)
where [x, y] is the vehicle position, $\beta$ is the slip angle at the center of gravity.
According to the aforementioned control module and kinematic module, the environment module can update the scenario information for the next time step according to CAV’s semantic decision. During this approach, the vehicle positions are calculated using precise mathematical formulas, thereby avoiding the potential uncertainties and errors that might arise from LLM directly controlling the vehicle.
III-C Reasoning Module
In this subsection, we establish an integrated reasoning module that progresses from state sharing to intent sharing, negotiation, and finally, decision. This module operates in a chain of thought manner, ensuring a smooth transition for CAVs from environmental perception to interaction and negotiation, and ultimately to decision-making.
The reasoning module first extracts information about surrounding vehicles from the environment to create a scene description. It then organizes the states of the vehicles into conflict pairs, forming conflict descriptions. To ensure consistency in vehicle decision-making during conflicts and to avoid collisions, we developed an LLM-based conflict coordinator. This coordinator integrates the current conflict descriptions with traffic rules to determine the order of priority for each conflict group. Finally, each CAV makes its decisions based on the conflict coordinator’s recommendations and its own scenario description.
III-C1 State-perception
The state-perception function is responsible for acquiring and processing information about the current environment of the CAV, including dynamic data such as lane information and vehicle information. The designed state-perception function aligns with the first level of CDA in the SAE J3216 standard, which is state sharing. CAVs are allowed to exchange information with others, therefore paving the way for subsequent higher levels of CDA. Through comprehensive analysis of the above information, the state-perception function can construct a complete and accurate driving environment recognition, which provides a reliable basis for subsequent reasoning.
Specifically, lane information is divided into three categories: the ego lane, adjacent lanes, and conflict lanes. The ego lane is the lane where the ego vehicle is currently driving in, adjacent lanes are the lanes on the left and right sides of the ego lane if they exist, and conflict lanes are the lanes that intersect with the ego lane. Similarly, vehicle information can be grouped based on their relationship with the ego vehicle into leading vehicles, rearing vehicles, surrounding vehicles, and conflict vehicles. During the state-perception process, lane information and vehicle information are combined to create an overview of the ego vehicle’s surrounding environment. Since vehicles in different lanes can affect the ego vehicle differently, a three-level action safety assessment is designed to ensure safe driving, which will be detailed in the decision subsection.
III-C2 Intent-sharing
The intent-sharing function, which conveys the driving intentions of a vehicle to other CAVs, is a key advantage of cooperative driving. From macro to micro, driving intention mainly includes the sharing of expected lanes and expected speed. Through intent-sharing, other vehicles can better understand the ego vehicle’s intentions, allowing them to make decisions while avoiding conflicts as much as possible.
State perception and intent sharing are combined as the scene descriptions. This description is based on the lane and includes all objects that could affect the ego vehicle, such as adjacent vehicles and conflicting vehicles. Therefore, it can be migrated to multiple scenarios, like intersections, highways, and merging areas, significantly enhancing the versatility of the proposed cooperative driving framework.
III-C3 Negotiation
CAVs possess broader perception capabilities and more powerful computing capabilities, which theoretically enable faster and more accurate recognition of the intentions of other vehicles. However, according to California’s Department of Motor Vehicles [25], over 31% of accidents are caused by CAV misjudging the intentions of other traffic participants, highlighting that traditional decision-making methods still have significant shortcomings in processing high-dimensional information such as intentions. Therefore, a conflict coordinator is designed to resolve conflicts and achieve level-3 CDA, which is known as agreement-seeking cooperation.
The conflict coordinator identifies all potential conflicts in the current environment and assesses the severity of each conflict based on the current state of both vehicles involved. To quantify the degree of conflict severity, the time difference to the conflict point as the surrogate indicator, defined as:
$\Delta TTCP$ $=\begin{vmatrix}TTCP_{i}-TTCP_{j}\end{vmatrix}=\begin{vmatrix}\frac{d_{i}}{v_{i}}-\frac{d_{j}}{v_{j}}\end{vmatrix}$ (7)
=$=\begin{cases}\Delta TTCP\leq 2s,\quad\text{Serious Danger}\\ 2s<\Delta TTCP\leq 5s,\quad\text{General Danger}\\ 5s<\Delta TTCP\leq 8s,\quad\text{Slight Danger}\\ \Delta TTCP\geq 8s,\quad\text{No Danger}\end{cases}$
where $TTCP$ is the time to conflict point based on the vehicle’s current distance to conflict point $d$ and speed. Among them, when $\Delta TTCP$ is less than 2s, it signifies a serious conflict, requiring at least one of them must take an emergency brake. If $\Delta TTCP$ falls between 2s and 5s, it is considered that there is a general conflict between the two vehicles, and at least one vehicle should slow down and yield. When the $\Delta TTCP$ is between 5s and 8s, the vehicles are considered to have a slight collision, and neither vehicle should accelerate simultaneously. When $\Delta TTCP$ is greater than 8s, it is considered safe that the conflict will not cause a collision.
The conflict coordinator determines the passing sequence of vehicles in each conflict pair based on the severity of the conflict. In this process, traffic rules and social norms in the driving process are simultaneously considered by the conflict coordinator. For example, as shown in Fig. 1, according to traffic rules, turning vehicles should yield to going-straight vehicles. Hence, after negotiation, the conflict coordinator determines that $\text{CAV}_{2}$ should yield in this pair of conflicts. The negotiation result and reasons are sent to the decision function to make the final decision. However, it is important to note that the negotiation result is advisory, and the final decision also depends on other factors in the ego vehicle’s surrounding environment.
III-C4 Decision
The decision module synthesizes the information from the three aforementioned modules to generate the final decision. Considering that the strength of LLMs lies in their reasoning capabilities based on world knowledge rather than numerical computation, we design the decision actions as discrete semantic decisions. The action space $A$ is designed as a set of high-level meta-actions. These meta-actions are then mapped to specific reference acceleration $a_{r}$ and reference heading $\phi_{r}$ values through a numerically based control model. Moreover, the vehicle’s next state is determined based on Eq. 5 and Eq. 6 to ensure safe and stable control.
The decision function mainly includes two tasks: scenario prompt generation and action safety assessment. Scenario prompt generation involves integrating the scene descriptions produced by state perception and intent sharing, as well as the negotiation results from the conflict coordinator, to summarize the key information in the surrounding driving environment. This integration serves as the prompt for LLM decision-making. Furthermore, compared with traditional decision-making methods, the safety of LLM decisions is often difficult to guarantee. Therefore, a three-layer action safety assessment method is proposed to determine whether action would pose safety risks in the current driving environment.
Based on lane classification, the first layer check determines whether the action will cause collisions with vehicles in the same lane, such as excessive acceleration leading to rear-end collisions or excessive deceleration causing abrupt braking by the rearing vehicle. Secondly, when the ego vehicle intends to take a left or right lane change, check whether the lane-changing will collide with other vehicles in the adjacent lane. Finally, check whether the action will lead to an intensification of the conflict with the vehicle in the conflict lane, especially avoiding serious and general dangers. During action safety assessment, any action found to have safety issues in the above checks is removed from the alternative action set. Therefore, by systematically assessing the feasibility of each action, the safety of LLM’s final decision is significantly improved. Fig. 3 summarizes the composition of LLM’s prompt in the inference process with CoT.
Figure 3: Prompt for LLM Reasoning Module.
III-D Memory Module
Enhancing continuous learning capabilities in autonomous systems has always been a significant challenge. Novice drivers accumulate experience through continuous driving practice, evaluate the effects of different behaviors, and learn from them to improve their driving skills. Drawing on this mechanism, a memory module is introduced to enable CAVs to learn from past experiences and utilize this knowledge to future interactions. This process is also referred to as Retrieve-Augmented Generation (RAG).
RAG endows LLMs with the ability to access specific knowledge databases within a domain or organization. This capability allows for economically efficient improvement of LLM outputs without requiring model retraining, ensuring relevance, accuracy, and practicality in addressing specific domain problems. Specifically, the designed memory module contains two primary functions: memory augment and memory retrieval.
III-D1 Memory augment
The memory augment function evaluates the impact of CAV actions from the previous scenario to determine whether these actions have exacerbated conflicts. If a CAV’s behavior leads to increased danger, the system generates negative feedback, such as: ”Your action has intensified the conflict; similar actions should be avoided.” This feedback mechanism establishes a connection between scenarios, actions, and results, storing these mappings in the memory databases for future reference. Before each invocation of the LLM for reasoning, the most relevant memories are retrieved from the memory databases to augment the prompts, thereby avoiding the repetition of past mistakes.
III-D2 Memory retrieval
As the number of interactions increases, the memory databases will accumulate numerous past experiences. Inputting all memories as prompts would lead to redundancy, making it difficult for the CAV to extract key information during inference. To address this problem, a memory retrieval function is employed to extract the most relevant memories to the current scenario from the databases before using memory to guide reasoning.
Specifically, current scenario descriptions and conflict descriptions are converted into vector forms and use cosine similarity to rank memories in the library based on their relevance to the current scenario. The top-ranked memories are then selected as part of the prompts for CAV inference. These similar memories, referred to as few-shot experiences, are injected into the CAV’s reasoning module, enabling the CAV to learn from past mistakes.
The introduction of the memory module not only improves the CAV’s decision-making ability in complex environments but also imparts human-like continuous learning capabilities. By continually learning from past experiences, CAVs can better adapt to dynamic environments, enhancing driving safety, reducing traffic accidents, and increasing the reliability and practicality of their real-world applications.
Algorithm. 1 summarizes the main process of our framework. In the next section, we verify the effectiveness of our framework in different scenarios through simulation experiments.
Input: Vehicle states $\boldsymbol{S}_{t}$, CAV list $\boldsymbol{C}$
Output: Vehicle next states $\boldsymbol{S}_{t+1}$
// Reasoning Module
1 Initialize the decision buffer of each CAV $\boldsymbol{D}\leftarrow[]$;
2 Initialize the scenario description buffer of each CAV $\boldsymbol{sce}\leftarrow[]$;
3 foreach $i\in\boldsymbol{C}$ do
4       Generate the scenario description $sce_{i}$ of vehicle $i$ based on current states $\boldsymbol{S}_{t}$;
5       Add scenario description to buffer $\boldsymbol{sce}\leftarrow\boldsymbol{sce}.\text{Add}(sce_{i})$ ;
6
7 end foreach
8Generate the conflict description $\boldsymbol{con}$ with conflict coordinator based on scenario description $\boldsymbol{sce}$ ;
9 foreach $i\in\boldsymbol{C}$ do
10       Sort the conflict description related to the ego vehicle $con_{i}\leftarrow\boldsymbol{con}.\text{Sort}(i)$;
       // Memory Module
11       Retrieve the most relevant memories $m$ to the current scenario $sce_{i}$ and conflict $con_{i}$ ;
12       Generate the prompt of decision based on memories $m$, scenario description $sce_{i}$, and conflict description $con_{i}$ ;
13       Reasoning the final semantic decision $d_{i}$ ;
14       Add decision to buffer $\boldsymbol{D}\leftarrow\boldsymbol{D}.\text{Add}(d_{i})$ ;
15
16 end foreach
// Environment Module
17 Update vehicle next state $\boldsymbol{S}_{t+1}$ with decision buffer $\boldsymbol{D}$ based on Eq. 5 and Eq. 6;
18 Evaluate the impact of the decision and add to the memory database;
Algorithm 1 LLM-driven cooperative driving framework.
IV Experiments and Analysis
To verify whether the proposed CoDrivingLLM can effectively improve the interaction ability and learning ability of CAV, this section carries out verification from three aspects. Firstly, ablation experiments were conducted on the negotiation module in different scenarios. Second, we compared the performance of CoDrivingLLM in different scenarios under 0-shot, 2-shots and 5-shots experiences. Finally, we evaluated the safety and efficiency of CoDrivingLLM against other cooperative driving methods, including optimization-based, rule-based, and machine learning-based methods.
IV-A Experiment Settings
Figure 4: Three scenarios we utilize to train and test our method, (a) scenario 1, highway environment; (b) scenario 2, merge environment; (3) scenario 3, intersection environment.
Simulation Environment. We develop our environment module based on the highway-env[26], an open-sourced and flexible simulation platform for autonomous driving. Three scenarios are designed to conduct experiments, as shown in Fig.4, including a four-way highway scenario, merge scenario, and single-lane unsignalized intersection. The settings of context traffic flow and HDVs follow the instructions of the environment module of our framework.
Implementation details. The GPT-4o mini is utilized our base LLM model to conduct high-level logical thinking and judgment. Four CAVs are controlled in each simulated environment. Furthermore, all scenarios with different settings are repeated 20 times with different random seeds to attain the final results. The success rate is used as the indicator to evaluate the performance of all methods. In one case, if all CAVs finish their driving tasks safely and arrive at their destination, this case is successful. Success rate means the ratio of the number of successful cases to the total number of cases.
IV-B Ablation Study on Negotiation
Lack of interaction ability is an important reason for the current CAV’s struggle on the open road. Cooperative driving, through the sharing of states and intentions, can provide additional information to help conflicting parties align their decision-making. To achieve this, we introduce a negotiation module into the reasoning process. By establishing an LLM-driven conflict coordinator, the traffic rules and scenario description are integrated to produce a suggested traffic sequence to assist the final decision, thus improving the interaction ability. To verify the effectiveness of the conflict coordinator in improving interaction, we conduct ablation experiments comparing the success rate of CoDrivingLLM in various scenarios with and without the negotiation module. In addition, we also select an example case to analyze how the negotiation function can improve safety.
Figure 5: The ablation experiment on negotiation module.
As depicted in Fig.5, in the highway scenario, the presence or absence of the negotiation module does not affect the performance of cooperative driving, as our method consistently achieves a 100% success rate. We attribute this to the lack of significant interaction conflicts in highway scenarios, where the negotiation module does not have an opportunity to intervene effectively. However, in the merge and intersection scenarios, the negotiation module significantly improves the success rate of decision-making. In these two types of scenarios, inevitable conflicts arise between vehicles, and collisions can only be avoided if their decision-making is aligned. Specifically, in the merge scenario, the success rate of CAV decisions increases from 33% to 75%, and in the intersection scenario, it rises from 15% to 78%.
The negotiation module has the most substantial impact in the intersection scenario, where the complexity of interactions and the density of traffic conflicts are much higher than in the other two scenarios. This demonstrates that our negotiation module effectively handles complex traffic conflicts and interaction scenarios, making it a critical component of the overall cooperative driving framework.
Figure 6: The cooperative driving cases at unsignalized intersection, (a) case from CoDrivingLLM without negotiation module, (b) case from CoDrivingLLM.
To better visualize the role of the negotiation module in the actual reasoning process, we selected a case with the same initial vehicle states in an intersection scenario to compare the different behaviors of vehicles with and without the negotiation module. Interaction cases videos and raw data of other scenarios can be found here [^1] https://fangshiyuu.github.io/CoDrivingLLM/.
In Fig.6(a), CoDrivingLLM is not equipped with the negotiation module. At 2 seconds, the decision module fails to effectively predict potential interaction conflicts, and the lack of necessary coordination between CAVs leads to a situation at 4 seconds where the decision module cannot generate a safe and effective passing strategy. As a result, $\text{CAV}_{4}$ continues to accelerate, ultimately colliding with $\text{CAV}_{5}$. We can observe that in such complex conflict scenarios, even when LLMs are provided with clear scene information and safety assessment methods, they still cannot eliminate all interaction conflicts and traffic risks, which is unacceptable in real-world traffic.
In contrast, as shown in Fig.6(b), CoDrivingLLM with the negotiation module effectively resolves this issue. In the same scenario, at 2 seconds, the negotiation module successfully identifies the potential movement conflict and, based on preset rules, suggests a recommended passing order, instructing $\text{CAV}_{4}$ to yield to $\text{CAV}_{5}$. This decision is ultimately adopted by the LLM in the decision module, leading $\text{CAV}_{4}$ to slow down at 4 seconds, allowing $\text{CAV}_{5}$ to pass safely, and successfully resolving the conflict.
This case demonstrates that our negotiation module effectively mitigates traffic conflicts and enhances the performance of the CoDrivingLLM method.
IV-C Comparison with Memory Module
Figure 7: The performance evaluation of memory module with different shots experiences.
The memory module helps CAVs learn from past experiences to avoid repeating mistakes. To validate the effectiveness of the memory module, we compare the performance of our method in different scenarios under 0-shot, 2-shots, and 5-shots experiences. In this context, 0-shot means the memory module is not used, while 2-shots and 5-shots refer to the introduction of the two and five most similar past experiences to the current scenario as prompts before reasoning.
As shown in Fig.7, in the highway scenario, where vehicle conflicts and interaction challenges are minimal, a 100% safety rate is achieved even without the memory module, and this success rate remains consistent when 2-shots and 5-shots experiences are added. However, in the merge and intersection scenarios, increasing the memory module from 0-shot to 2-shots significantly improves the decision-making success rate, with the merge scenario rising from 78% to 90% and the intersection scenario from 75% to 85%.
We do note that when the memory module is further increased from 2-shots to 5-shots, there is a slight decline in performance in both scenarios. We believe this indicates that longer memory does not necessarily enhance LLM decision-making. An excess of repetitive or low-value memories may occupy too much of the LLM’s context window, reducing its ability to focus on other relevant perceptual information and potentially leading to a degree of negative impact.
Figure 8: The cooperative driving cases at unsignalized intersection, (a) case from CoDrivingLLM without memory module, (b) case from CoDrivingLLM.
We also use an intersection scenario case to demonstrate the role of the memory module. Fig.8(a) shows a typical case of decision failure. In this instance, CoDrivingLLM is not equipped with the memory module. At 4 seconds, although the decision module, through the negotiation module, identifies a potential rear-end conflict between $\text{CAV}_{7}$ and $\text{CAV}_{4}$ and attempts to resolve it by adjusting actions, $\text{CAV}_{4}$ accelerates while $\text{CAV}_{7}$ decelerates. However, due to $\text{CAV}_{7}$’s high initial speed, it collides with $\text{CAV}_{4}$ before it can fully decelerate. This case illustrates that in certain special scenarios, cooperative decision-making can still fail even with the negotiation module in place.
In contrast, Fig.8(b) demonstrates the decision-making performance of CoDrivingLLM after utilizing the memory module. By retrieving memories, CoDrivingLLM identifies a similar scenario where a decision failed and conducts a few-shot experience learning session. At 4 seconds, it outputs a different decision strategy, instructing $\text{CAV}_{7}$ to accelerate and pass while $\text{CAV}_{4}$ slows down to yield. This decision is the opposite of the one in Fig.8(a) and effectively resolves the conflict between the two CAVs, ultimately allowing all CAVs to reach their destinations safely.
This case effectively showcases the learning utility provided by the memory module to CoDrivingLLM.
IV-D Comparison with Other Methods
Considering that not all cooperative driving methods are suitable for multiple scenario, and unsignalized intersections are often where accidents occur most frequently [27]. Therefore, we selected the three typical cooperative driving methods mentioned earlier to test our approach under unsignalized intersection scenario, including optimization-based method (iDFST)[9], rule-based method (Cooperitave game)[28], learning-based method (MADQN)[29].
TABLE I: Comparison under various methods
Methods Success rate Post Encroachment Time Travel Velocity
Average Max Min Average Max Min
iDFST 85% 15.1 47.6 1.2 4.1 7.9 0.2
Cooperative game 55% 5.7 18.6 0.8 5.7 7.9 0.9
MADQN 20% 3.7 9 $\boldsymbol{1.8*}$ $\boldsymbol{6.1*}$ 6.5 5.4
CoDrivingLLM (Ours) 90% 10.3 44 $\boldsymbol{1.8*}$ 5.7 8 1
IV-D1 Overall Performance
Firstly, we summarized the success rate of each methods. According to the data presented in Table. I, CoDrivingLLM outperforms the other methods, achieving a success rate of 90%. The iDFST and Cooperative Game methods lag slightly behind, both achieving a success rate of 85%. In contrast, MADQN displays the poorest generalization and performance, with a success rate of merely 20%.
Figure 9: PET evaluation results of different methods.
IV-D2 Safety Evaluation
We have conducted a comprehensive safety analysis of various methods, utilizing the Post-Encroachment Time (PET) metric, a widely recognized safety parameter in traffic engineering. This metric effectively quantifies the safety and interaction intensity of vehicles in intricate traffic scenarios. According to the results presented in Fig. 9 and Table. I, both the iDFST approach and CoDrivingLLM exhibit commendable performance, with mean PET values of 15.1 seconds and 10.3 seconds, respectively. Conversely, MADQN and the Cooperative method demonstrate PET values of 5.7 seconds and 3.7 seconds, respectively, indicating a heightened risk in decision-making interactions within the traffic environment. In addition, a lower PET value signifies an elevated safety hazard. Normally, a PET value less than 1.5s signifies severe conflicts that should be avoided. According to Fig. 9, The PET distribution of the proposed CoDrivingLLM is predominantly above 1.5s. Although the PET distribution for MADQN shows no values below 1.5s as well, considering that its success rate is only 20%, there are a large number of collisions. Therefore, CoDrivingLLM performs best in terms of safety.
IV-D3 Efficiency Evaluation
We also evaluate the efficiency of all methods using travel velocity. As shown in Table. I, MADQN achieves the highest average speed at 6.1 m/s during the tests. However, this efficiency comes at the cost of significantly compromising safety and overall performance. The iDFST method performs well in terms of safety and success rate, it exhibits poor driving efficiency, with excessive low-speed driving resulting in an average speed of only 4.1 m/s. Although the rule-based approach performs well in terms of safety, its low efficiency has consistently been a common issue. In contrast, CoDrivingLLM strikes a balance between safety and efficiency, maintaining a high average speed of 5.7 m/s while ensuring safety.
In summary, CoDrivingLLM effectively balances safety and efficiency, achieving the best overall performance, which clearly demonstrates its superiority.
Figure 10: The average speed of different methods.
V Conclusion
Cooperative driving technology effectively enhances the quality of cooperative decision-making for CAVs. Leveraging the rapid development of LLMs in recent years and their demonstrated strong logical reasoning and multi-task generalization capabilities, we propose an interactive, learnable LLM-driven cooperative decision-making framework. This framework includes three modules: the environment module, reasoning module, and memory module. Within this framework, we integrate four stages of cooperative decision-making, including state sharing, intent sharing, negotiation, and action based on different levels of CDA. Each CAV performs distributed high-level logical reasoning using LLMs, based on its own observations and the shared state and intent information from other vehicles, with centralized conflict resolution managed through a conflict coordinator within the framework. Additionally, we design a memory module to store decision-making processes and outcomes, enabling CAVs to achieve continuous learning and self-evolution through RAG technology. Our method is tested across multiple cooperative scenarios, and the results demonstrate that our approach significantly outperforms other rule-based, optimization-based, and machine-learning methods in terms of success rate, safety, efficency.
In the future, we aim to broaden the application of this framework to encompass a wider range of scenarios. We plan to undertake large-scale cooperative decision-making experiments involving CAVs, and we aspire to deploy this framework on actual experimental vehicles in real-world settings. Additionally, we intend to refine the memory module and evaluate various LLM base models to bolster the practical applicability and performance of our approach.